# Quantum mechanics provides the physical basis of teleological evolutions

### v4

@@ 2.7 About the teleological character of the evolutions of the living
Note that a teleological character of the evolutions of the living on the Earth could be further explored by investigating, either theoretically or by means of toy models, the character of the operation of “tracing out all the rest”. We should consider the unitary quantum evolution of a system endowed with quantum speedup, standing for the evolution of the universe toward life, and derive from it the almost autonomous evolution of a subsystem of it, standing for the evolution of life on Earth, by tracing out all the rest. We should look for a way of obtaining an evolution of the subsystem that is teleological (thus quantum) in the longer term and (about) classical in the shorter term. If this were possible, it would go in favor of the quantum cosmological reading of the Anthropic Principle we are dealing with. By the way, this could also be another perspective under which to address the problem of unifying quantum and classical theory.
@@
Let us end the present section by recalling what, to our knowledge, are the key steps that led to the discovery of the quantum computational speedup. According to the present theory, this discovery would be even more fundamental than currently thought. The first step was showing that computation is possible in the quantum world and introducing the fundamental notion of quantum bit, by Finkelstein [21] in 1969; that was also the first time that quantum physics moved to the world of the complex, indeed to a cellular automaton model of the universe—this was in the context of his well known attempt to unify relativity and quantum theory. Then, there was the introduction of the notion of logically reversible computation by Bennett [22] in 1973, of course, a prerequisite for the building up of computations in unitary quantum evolutions. The discovery of the superior computational efficiency of the quantum world with respect to that of the classical world in 1982 is due to Feynman [23]. Still in 1982, there has been the introduction of the notion of conservative logic by Fredkin and Toffoli [24]: with it, we have the logically reversible gates needed to perform quantum computations. The seminal quantum algorithm that delivers a quantum speedup was conceived by Deutsch [25] in 1985.
Note the importance of the role that the notion of reversibility played in the development of quantum computation. The present work could be put on the same path. It could be said that the present time-symmetric form of causality exploits the objectively ideal time-reversibility of unitary quantum evolutions in the deepest possible way.
3 Precedents
We have seen that applying the quantum correlation time-symmetrization rule to quantum nonlocality provides a local explanation of it. Now, we recall a historical precedent of this.
It is a work by Costa de Beauregard [26, 27] that also provides a retrocausal explanation of quantum nonlocality. In it, the outcome of the first performed measurement—of the observable of one of the two space separate subsystems—locally and unitarily goes backward in time until the time the two subsystems were not yet space-separate, where it locally tells the other subsystem the state to assume in view of the future measurement. Naturally, we are not dealing with the present time-symmetric form of retrocausality, but the concept is similar.
Speaking of precedents, of course, the present work is inspired by time-symmetric quantum mechanics [28, 29]. Furthermore, there have been and there are studies on the possible violations of the causal principle by quantum mechanics. However, none of them concerns the quantum computational speedup. Their possible relation to the present work, as things are now, would not be evident.
@@ 4 Conclusions
The present work goes against a long-standing de facto no-retrocausality tradition of physics. More specifically, since causality could well go backward in time along a unitary quantum evolution that is both objectively and ideally time-reversible, we should say that, in the quantum world, the scientific tradition amounts to asserting the mutual exclusivity between the two possible time directions of causality.
Let us first address the reliability of the work, a particular issue because of its unconventionality.
That the classical logic description of the quantum speedup must resort to retrocausality should be obvious. Of course, it should be obvious the fact that always locating a ball hidden in one of four drawers by opening a single drawer is logically equivalent to always knowing in advance that the ball is in a pair of drawers. To the problem-solver, the only possible source of this information is the solution of the problem she will produce and measure in the future: hence retrocausality. We have seen how this immediately generalizes to the full fledged Grover algorithm and a most general form of quantum correlation.
At this point, we should address the legitimacy of applying classical logic to what the quantum algorithm does. Let us recall our two arguments in favor of this choice.
The first is that, only involving the measurements of commuting observables (the problem-setting and the solution of the problem), quantum algorithms (and what they do) should be subject to classical logic.
The other is that all the results obtained under the assumption that quantum algorithms can be described by classical logic are also obtained by time-symmetrizing their ordinary quantum description; this, in turn, is mandatory to satisfy the atemporal symmetry between the simultaneous measurements of two maximally entangled observables implicit in the ordinary quantum description of quantum algorithms.
We would think that the present retrocausal explanation of the quantum speedup is reliable enough to deserve a discussion despite its destabilizing consequences.
The flip side of the destabilizing character of the work is its innovative character. Let us distinguish between what should be certain and what should be plausible. What should be certain: (i) The work provides a unified quantitative explanation of the quantum speedups that is still missing from mainstream quantum mechanics. (ii) It provides a theoretical confirmation of the incompleteness of the quantum description of quantum nonlocality sustained by Einstein and others, also providing the local hidden variables envisaged by those authors to complete it. (iii) It highlights the teleological character of quantum algorithms; given the negative view that the teleological notion has in today’s science and the scope of this notion, this should be the most important result. What should be plausible: It shows that, under a quantum cosmological reading of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living.
In conclusion, the present work should show that there is another possible view of physical reality out there, outside the current view. We note that this might not only be a curiosity. A revamping of the abandoned notion of the relation between evolution and functional structure of the living might have a profound impact on today’s evolutionary biology and biology.
@@
Also note that table III diagram goes along with the fact that the unitary part of the quantum algorithm, $\hat{\mathbf{U}}_{1,2}$, takes the opening just one drawer, but does not imply it. We know this fact from outside Table III diagram: we know that this is what the unitary part of Grover algorithm takes. The diagram says that the problem-solver knows in advance that the ball is in a pair of drawers rather than four—bottom left corner of the diagram—and this could allow her to solve the problem by opening a single drawer—bottom line of the diagram of table III or the diagram of table IV. The advanced knowledge rule tells us the room for the quantum speedup but does not imply it.

### v7

@@ 3 Discussion
Note that a teleological character of the evolutions of the living on the Earth could be further explored by investigating, either theoretically or by means of toy models, the character that the operation of “tracing out all the rest” would acquire in the present context. We should consider the unitary quantum evolution of a system endowed with quantum speedup, standing for the evolution of the universe toward life, and derive from it, by tracing out all the rest, the almost autonomous evolution of a subsystem of it, standing for the evolution of life on Earth. We should look for a way of obtaining an evolution of the subsystem that is teleological (thus quantum) in the longer term and (about) classical in the shorter term. If this were possible, it would go in favor of the quantum cosmological reading of the Anthropic Principle that we have proposed. By the way, this could also be another perspective under which to address the problem of unifying quantum and classical theory, in this case under the present form of time-symmetric quantum mechanics.
The same might also hold for cosmology. After all, under the quantum cosmological assumption, also the physical part of the evolution of the universe toward life should be subject to the Fine-tuned Universe version of the Anthropic Principle. Also this part could be endowed with quantum speedup and thus be teleological in character. It might be interesting to confront this possibility with the present cosmological notions, naturally comprising their open problems.
We have already noted that the present notion of teleological evolution implies the quantum speedup of the latter. This would, of course, be a novelty with respect to the historical notion of teleological evolution. It might even better explain than Darwinian gradualism the extraordinary speed with which some evolution steps occurred.
@@ 4 Conclusions
The present work goes against a long-standing de facto no retrocausality tradition of physics. More specifically, since causality could well go backward in time along a unitary quantum evolution that is both objectively and ideally time-reversible, we should say that, in the quantum world, the scientific tradition amounts to asserting the mutual exclusivity between the two possible time directions of causality.
Let us first address the reliability of the present unconventional work.
The logic for which quantum algorithms should be teleological evolutions is simple. Always locating a ball hidden in one of four drawers by opening a single drawer, as the quantum algorithm does, is obviously equivalent to knowing in advance that the ball is, in a mutually exclusive way, in one of the three pairs of drawers in which it must be. This, of course, implies knowing in advance, in a mutually exclusive way, one of the possible halves of the information about the solution one will produce and measure in the future and using this knowledge to produce the solution with fewer computation steps. This attraction from the future is indeed the teleological character of the quantum algorithm. It should be completely evident. We have seen that this immediately generalizes to any number of drawers and most general forms of quantum correlation.
The above is also confirmed by the fact that the time-symmetrization of the ordinary quantum description of the quantum algorithms says the same thing and is mandatory to satisfy the atemporal symmetry implicit in the latter description.
Summarizing, the teleological character of quantum algorithm should be evident, and enough to dispel the belief that teleological evolutions have no physical basis.
@@
Also note that table III diagram goes along with the fact that the unitary part of the quantum algorithm, $\hat{\mathbf{U}}_{1,2}$, takes the opening of just one drawer, but does not imply it. We know this fact from outside Table III diagram: we know that this is what the unitary part of Grover algorithm takes. The diagram says that the problem-solver knows in advance that the ball is in a pair of drawers rather than four—bottom left corner of the diagram—and this could allow her to solve the problem by opening a single drawer—bottom line of the diagram of table III or the diagram of table IV. The advanced knowledge rule tells us the room for the quantum speedup but does not imply it.